OpenStudy (anonymous):

Find the probability that a point chosen randomly inside the triangle is within the square. Round to the nearest hundredth.

OpenStudy (anonymous):



OpenStudy (welshfella):

required probability = area of the square / area of the triangle

OpenStudy (welshfella):

= 4^2 / (1/2 * 8 * 10)

OpenStudy (anonymous):

I don't understand!!

OpenStudy (welshfella):

the square is inside the triangle so chances of being in the square is area of sqaure divided by the area of then whole thing ( the triangle)

OpenStudy (anonymous):

so is it 16/40= 4/10= 2/5